OpenStudy (anonymous):

on saturday tom worked three less hours than on friday. If he worked 10 hours in the two days, how many did he work on friday?

OpenStudy (anonymous):

There is no 1 1/2 for an answer...

OpenStudy (jiteshmeghwal9):

Let the no. of hours he worked in Friday be \(x\). then, so on Saturday, he works=(x-3) The sum of the hours is 10. so, x+(x-3)=10

OpenStudy (jiteshmeghwal9):

solve for x & u will gt ur answer :)

OpenStudy (anonymous):

6 1/2?

OpenStudy (anonymous):

s + f = 10 s = f - 3 now sub f - 3 in for s in the first equation f - 3 + f = 10 2f - 3 = 10 2f = 10 + 3 2f = 13 f = 13/2 = 6 1/2 now sub 6 1/2(or 6.5) in for f into the 1st equation s + f = 10 s + 6.5 = 10 s = 10 - 6.5 s = 3.5 check... s = f - 3 3.5 = 6.5 - 3 3.5 = 3.5 (correct) or... s + f = 10 3.5 + 6.5 = 10 10 = 10 (correct) On Friday, Tom worked 6.5 hours

OpenStudy (anonymous):

on a trip that took 10 hours, mark drove 2 fewer hours than mary. how many hours did mary drive?

OpenStudy (jiteshmeghwal9):

Mary's hours=\(x\) Mark's hours=\(x-2\) Since on the trip it took 10 hours. therefore, \(x+(x-2)=10\)